OpenStudy (anonymous):

Solve the equation by factoring. 4x2 - 7x = 2

OpenStudy (anonymous):

\[4x^2-7x=2\]

OpenStudy (lgbasallote):

first step is to subtract 2 from both sides

OpenStudy (anonymous):

ok then divide by 4

OpenStudy (anonymous):

If it is to the power of whatever you put ^

OpenStudy (anonymous):

for example x^2;)

OpenStudy (anonymous):

do i do formula \[b+-\sqrt{b^2-4ac}/2a\]

OpenStudy (anonymous):

does anyone know the answer

OpenStudy (anonymous):

ax^2+bx+c=0 4x^2-7x-2=0 Then plug in a b and c

OpenStudy (anonymous):

7+ the square root of (7^2-4*4*2) divided by 2*4

OpenStudy (anonymous):

and 7- the square root of (7^2-4*4*2) divided by 2*4

OpenStudy (anonymous):

which is

OpenStudy (anonymous):

Do both then plug them in

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Since we want to solve this by factoring, one side of the equation must be zero. That's a key step. \[4x^2-7x=2\] \[4x^2-7x-2=2-2\] Subtract both sides by 2 to make the right side 0. \[4x^2-7x-2=0\]

OpenStudy (anonymous):

yes i got that! then what

OpenStudy (anonymous):

Next step is to factor out the trinomial on the left side.

OpenStudy (anonymous):

To do that, think of two numbers such that the product is -8 and sum is -7.

OpenStudy (anonymous):

The correct combination is -8 and 1. Since (-8)(1)=-8 and (-8)+(1)=-7.

OpenStudy (anonymous):

Now we will replace the middle term of the trinomial by the two numbers we found. \[4x^2-7x-2=0\] \[4x^2-8x+1x-2=0\]

OpenStudy (anonymous):

Grouping them into two... \[(4x^2-8x)+(x-2)=0\] Factor out 4x on the first parenthesis, and factor out 1 on the second parenthesis... \[4x(x-2)+1(x-2)=0\]

OpenStudy (anonymous):

Notice that we can do one final factorization by pulling out the common factor (x-2). This gives us... \[(x-2)(4x+1)=0\]

OpenStudy (anonymous):

One last step. Set each factor equal to zero and solve for x. x-2= 0 x=2 4x+1=0 4x=-1 x=-1/4 So our final answers are x = 2 and x=-1/4. Hope this helps.

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (anonymous):

You're welcome. Thank you.